# ReZero: Boosting MCTS-based Algorithms by Backward-view and Entire-buffer Reanalyze

### v3

@@ 3.1 MuZero
3.2 Stochastic Bandits
In the upcoming Section 4.1, we plan to provide an inspiration case study behind our method, which is rooted in the fundamentals of bandit algorithms. Before diving into that, we’ll start with an overview of the key concepts surrounding stochastic bandits.
A stochastic bandit $\nu$ with $K$ arms [26] can be formulated as $\nu=\{P_{i}|i\in{1,2,...,K}\}$, where $P_{i}$ denotes the reward distribution corresponding to the arm $i$. The expected reward of arm $i$ is denoted as $\mu_{i}(\nu)$. Given an algorithm $\pi$ interacting with the bandit $\nu$ for $n$ rounds, we define the regret $R_{n}$:
$R_{n}(\pi,\nu)=n\mu^{*}(\nu)-\mathbb{E}[\sum_{t=1}^{n}X_{t}]$ (2)
where $\mu^{*}(\nu)=\max_{a\in[K]}\mu_{a}(\nu)$, $X_{t}$ denotes the random reward at time $t$. The regret quantifies the loss of reward incurred by $\pi$ compared to the optimal policy on $\nu$ during $n$ rounds of interactions. In practice, we aim for $\pi$ to exhibit sub-linear regret growth across a category of bandits $\Omega$ , i.e.:
$\lim_{n\to\infty}\frac{R_{n}(\pi,\nu)}{n}=0,\ \ \ \text{for all $\nu\in\Omega$ }$ (3)
where sub-linear growth rate indicates $\pi$ is choosing the optimal arm almost all of the time with $n$ tends to infinity. UCB [8] is a classical bandit algorithm that achieves regret growth rate of $O(\log(n))$.
AlphaZero [9] considers the child nodes in tree search as bandit arms and applies a variant of the UCB in the tree search process:
@@ 4 Method
4.1 An Inspiring Case Study
As a classic algorithm in the realm of stochastic bandits, Upper Confidence Bound 1 (UCB1) [8] has been deployed to select nodes within a search tree.
Figure 1: The connection between MCTS and bandits. Left shows the tree search in a bandit-view. When action $A$ is selected, a reward $r_{A}$ will be returned, where $r_{A}=\sum_{t=1}^{3}\gamma^{t-1}r_{t}^{a}+\gamma^{3}v_{a}$. For the root node, the traversal, evaluation and back-propagation occurring in the gray box can be approximated as sampling from an non-stationary distribution. Thus the root node can be seen as a non-stationary bandit. Right shows the one-armed bandit case. Once the reward $\mu_{A}$ is known, we can evaluate arm A using $\mu_{A}$, thereby eliminating the need to rely on subsequent tree search processes.
As explained in Figure 1, each node in the tree can be regarded as a multi-armed bandit Notice that in the context of stochastic bandits, the reward for each arm is sampled from a stationary probability distribution. However, in the process of tree search, the value obtained from selecting a node depends on subsequent search process. The UCT [27] algorithm models this process as sampling from a non-stationary distribution which converges. For the sake of analytical convenience in this paper, we approximate the root node as a stochastic bandit and consider each child node as an arm with a deterministic reward distribution. Based on this modeling, we can analyze the Eq. 4 employed in AlphaZero [9] and derive the following proposal:
Proposal 1
Denote the prior for arm $i$ as $P_{i}$. Without loss of generality, suppose $\mu^{*}=\mu_{1}$. Assume that the reward distribution of each arm $i$ is $1$-subgaussian and $P_{1}>\mu_{1}$, then the regret $R_{n}$ of using Eq. 4 to choose arm satisfies
$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (5)
The assumption that $P_{1}>\mu_{1}$ is made purely for the sake of formality; in fact, Appendix A indicates that, as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, a similar upper bound on regret can always be obtained. This elucidates that selecting actions according to Eq. 4 can yield a sublinear growth of regret, indicating that the final visit distribution will be concentrated on the optimal arm (see analysis for Lemma A.1 in Appendix A). This also confirms the rationality of selecting the action with the largest number of visits. The inspiration of our enhanced method comes from the one-armed bandit model: if the expected reward of a particular arm is known as prior, the necessity for exploratory trials on this arm is obviated, thereby reducing the regret upper bound. To concretize this insight, we can extend Eq. 4 and deduce the following proposition:
Proposal 2
For a bandit which satisfies the assumptions in Proposal 1, the regret upper bound can be reduced if we use the known expectation value of an arm to replace its score in Eq. 4. Specifically, if the known arm is the optimal arm with value $\mu_{1}$, then the regret $R_{n}$ satisfies:
$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(2+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (6)
In other case, with the known value $\mu_{l}$ of a sub-optimal arm $l$, the regret $R_{n}$ satisfies:
$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})+\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}})$ (7)
Proposal 2 demonstrates that, given the known expected reward of a suboptimal arm, the asymptotic upper bound of regret can be further reduced, thereby concentrating the visit distribution more heavily on the optimal arm. Meanwhile, as shown in Figure 1, when the expectation $\mu_{A}$ is known, we no longer need to evaluate arm A through real-time search results. Instead, we can directly substitute $\mu_{A}$ for the score in Eq. 4, thus saving the time spent on traverse and back-propagation processes. This inspires us to evaluate a particular child node using existing information, thereby eliminating the exploration of the corresponding sub-tree and accelerating the search process. See Appendix A for detailed proof and analysis of the two proposals.
@@ 4.2 Backward-view Reanalyze
Inspired by the aforementioned motivation, we aspire to obtain the reward expectation of a child node in advance to assess it. However, the true expectation is always unknown. Consequently, we resort to using the root value obtained from MCTS as an approximate substitute. Specifically, for two adjacent time steps $S_{0}^{1}$ and $S_{0}^{0}$ in Figure 2, when searching for state $S_{0}^{1}$, the root node corresponds to a child node of $S_{0}^{0}$. Therefore, the root value obtained can be utilized to assist the search for $S_{0}^{0}$. However, there is a temporal contradiction. $S_{0}^{1}$ is the successor state of $S_{0}^{0}$, yet we need to complete the search for $S_{0}^{1}$ first. Fortunately, this is possible during the reanalyze process.
Algorithm 1 shows the specific design with Python-like code. During reanalyze, since the trajectories was already collected in the replay buffer, we can perform tree search in a backward-view, which searches the trajectory in a reverse order. Figure 2 illustrates a batch containing $n+1$ trajectories, each of length $k+1$. We denote $S_{l}^{t}$ as the $t$-th state in the $l$-th trajectory. We first conduct a search for all $S^{k}$s, followed by a search for all $S^{k-1}$s, and so on. After we search on state $S_{l}^{t+1}$  [^1] To facilitate the explanation, we have omitted details such as the latent space and do not make a deliberate distinction between states and nodes., the root value $m_{l}^{t+1}$ is obtained. When engaging in search on $S_{l}^{t}$, we assign the value of $S_{l}^{t+1}$ to the fixed value $m_{l}^{t+1}$. During traverse in the tree, we select the action $a_{root}$ for root node $S_{l}^{t}$ with the following equation:
@@
where $a$ refers to the action associated with a child node, $a_{l}^{t}$ is the action corresponding to $S_{l}^{t+1}$, and $r_{l}^{t}$ signifies the reward predicted by the dynamic model. If an action distinct from $a_{l}^{t}$ is selected, the simulation continues its traversal in accordance with the original setting [10]. If action $a_{l}^{t}$ is selected, this simulation is terminated immediately. Since the time used to search for node $S^{t+1}_{l}$ is saved, this enhanced search process is faster than the original version. Notice that the analysis in Proposal 2 assumes that we know the expectation value. However, in our design, we use the root value as an approximate estimation. Still based on the stochastic bandit assumption, the following theorem analyzes the impact of the error between these two values:
Theorem 1
For a bandit which satisfies assumptions in Proposal 1 and a sample result for an arm, the regret upper bound can be reduced if we use the way of Equation 8 to choose arm. Specifically, if the sampled arm is the optimal arm with $n$ times sample mean $\hat{\mu}_{1n}$, then the regret $R_{n}$ satisfies:
$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(2+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+n\exp(-\frac{n\varepsilon^{2}}{2}))$ (10)
In other case, with the $n$ times sample mean $\hat{\mu}_{ln}$ of a sub-optimal arm $l$, the regret $R_{n}$ satisfies:
$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (11)
$+$ $\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}}+n\exp(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{2}))$ (12)
This theorem demonstrates that using the sampled mean as a substitute for the expectation value can also reduce the asymptotic upper bound of regret and maintain sub linear growth rate. This can be understood as follows: although there is an error between the sampled mean and the expectation value, since we use the average of $n$ samples (as the root value comes from a tree search which has $n$ simulations), this sampled mean will increasingly approximate the expectation value as $n$ grows, thereby preventing a linear increase. Refer to Appendix A for the proof.
@@ Appendix A Proof Materials
Regret decomposition lemma:Denote $X_{t}$ as the reward at time $t$, $A_{t}$ as the arm chosen at time $t$, and $S_{n}=\sum_{t}X_{t}$, then the regret can be decomposed by regret caused by every sub-optimal arm:
$R_{n}=\sum_{i}E[T_{i}]\Delta_{i}$ (13)
Proof.
Note that
$R_{n}$ $=n\mu^{*}-E[S_{n}]$ (14)
$=\sum_{i}\sum_{t=1}^{n}E[(\mu^{*}-X_{t})\mathbb{I}\{A_{t}=i\}]$ (15)
and
$E[(\mu^{*}-X_{t})\mathbb{I}\{A_{t}=i\}|A_{t}]$ (16)
$=\mathbb{I}\{A_{t}=i\}E[\mu^{*}-X_{t}|A_{t}]$ (17)
$=\mathbb{I}\{A_{t}=i\}(\mu^{*}-\mu_{A_{t}})$ (18)
$=\mathbb{I}\{A_{t}=i\}(\mu^{*}-\mu_{i})$ (19)
$=\mathbb{I}\{A_{t}=i\}\Delta_{i}$ (20)
The proof can be completed by combining the two equations and using the definition of $T_{i}$. ∎
This lemma demonstrates that for a certain bandit, if the regret grows sub linearly, then the expected times each sub optimal arm is chosen also grows sub linearly. Consequently, as $n$ increases, the proportion of times the sub optimal arm is selected converges to 00. In other words, the visit count distribution concentrates on the optimal arm as $n$ grows.
Lemma A.2.
Let $X_{1},...,X_{n}$ be a sequence of independent 1-subgaussian random variables, $\hat{\mu_{t}}=\frac{1}{t}\sum^{t}_{s=1}X_{s}$, $\varepsilon>0$, $a>0$ and
@@
, then it holds that $E[\kappa]\leq 1+\frac{2\sqrt{a}}{\varepsilon}+\frac{2}{\varepsilon^{2}}$.
@@
$E[\kappa]$ $\leq u+\sum_{t=\lceil u\rceil}^{n}\mathbb{P}(\hat{\mu}_{t}\sqrt{\frac{a}{t^{2}}}\geq\varepsilon)$ (23)
$\leq u+\sum_{t=\lceil u\rceil}^{n}\exp(-\frac{t(\varepsilon-\sqrt{\frac{a}{t^{2}}})^{2}}{2})$ (24)
@@
$=1+u+\frac{2}{\varepsilon^{2}}=1+\frac{2\sqrt{a}}{\varepsilon}+\frac{2}{\varepsilon^{2}}$ (27)
Proof for Proposal 1:
@@
Denote $T_{i}(k)$ as the times that arm $i$ has been chosen until time $k$, $\hat{\mu}_{is}$ as the empirical mean based on the first $s$ samples from distribution $P_{i}$, and $\hat{\mu}_{i}(k)=\hat{\mu}_{iT_{i}(k)}$. Then we have:
@@
$\leq 1+\sum_{t=2}^{n}\sum_{s=1}^{t-1}\exp(-\frac{s(\varepsilon+P_{1}\frac{\sqrt{t-1}}{1+s})^{2}}{2})$ (34)
$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{2}\varepsilon P_{1}\sqrt{t-1})\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{2})$ (35)
$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{2}\varepsilon P_{1}\sqrt{t-1})\frac{2}{\varepsilon^{2}}$ (36)
$\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (37)
Notes: In 32, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00. Thus, we can discard it. If this assumption doesn’t hold, we can choose to accumulate $t$ starting from a larger $t_{0}$(which satisfies $P_{1}\sqrt{t_{0}-1}>\mu_{1}$ as mentioned in the article). Starting the summation from such a $t_{0}$ ensures all terms of $s=0$ can still be discarded, and all add terms that $t\leq t_{0}$ can be bounded to $1$. This only changes the constant term and won’t affect the growth rate of regret. From 34 to 35, we just need to expand the quadratic term and do some simple inequality scaling. From 35 to 36, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{2})$ is a geometric sequence and scale it to $\frac{2}{\varepsilon^{2}}$. From 36 to 37, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
for 30, we have
@@
[…]
@@
$E[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (43)
combined with Lemma A.1, the proof is completed. ∎
[…]
Proof for Proposal 2:
Proof.
The proof can be obtained by making slight modifications to the proof of Proposal 1. In case that $\mu_{1}$ is already known,
$E[T_{i}(n)]=E\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}]$ (44)
$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\mu_{1}\leq\mu_{1}-\varepsilon\}]$ (45)
$=E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (47)
$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}$ (48)
In case that a sub-optimal value $\mu_{l}$ is known,
$E[T_{l}(n)]=E[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=l\}]$ (49)
$+E[\sum_{t=1}^{n}\mathbb{I}\{\mu_{l}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (51)
$=E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (52)
$\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (53)
and the bound of $T_{i}(n)$ for $i\neq l$ keeps unchanged. In both cases, combined with Lemma A.1, the proof is completed. ∎
From the proof, it can be seen that if the real value $\mu_{1}$ of the optimal arm is known in advance, the regret caused by the underestimation of the optimal arm can be eliminated. However, the part of the regret that is reduced is $O(1)$. If the real value $\mu_{i}$ of a sub-optimal arm $l$ is known in advance, the regret caused by continuous exploration of this sub-optimal arm can be eliminated. The part of the reduced regret is $O(\sqrt{n})$. Intuitively, it can be understood this way: If the value of a sub-optimal arm $l$ is known, then after sufficient exploration of the optimal arm, the upper confidence bound will converge to be higher than $\mu_{l}$ with probability $1$, thus making the regret caused by arm $l$ no longer increases with $n$. While if the value of the optimal arm is known, then as the number of simulations $n$ increases, there will always be a moment when the upper confidence bound of the sub-optimal arm $i$ is greater than $\mu_{1}$, so arm $i$ will still be repeatedly selected, leading to the regret caused by arm $i$ increases with $n$. However, in this case, since the optimal arm is chosen the most time as $n$ increases, we can save more time by early stopping the search.
Proof for Theorem 1:
Proof.
The proof can be obtained by making slight modifications to the proof of Proposal 2. In case we have $\hat{\mu}_{1n}=\frac{1}{n}\sum_{t=1}^{n}\hat{\mu}_{t}$ where $\hat{\mu}_{t}$ is sampled from the same reward distribution of arm $1$,
@@
$\leq n\exp{(-\frac{n\varepsilon^{2}}{2})}+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (57)
$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+n\exp{(-\frac{n\varepsilon^{2}}{2})}$ (58)
@@
$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{2})}$ (62)
$\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{2})}$ (63)
and the bound of $E[T_{i}(n)]$ for $i\neq l$ keeps unchanged.
In both cases, combined with Lemma A.1, the proof is completed. ∎

### v4

@@ 3.1 MuZero
3.2 Bandit-view Tree Search
A stochastic bandit has $K$ arms, and playing each arm means sampling a reward from the corresponding distribution. For a search tree in MCTS, the root node can be seen as a bandit, with each child node as an arm. The left side of Figure 1 illustrates this idea. However, as the policy is continuously improved during the search process, the reward distribution for the arms should change over time. Therefore, UCT [27] modeled the root node as a non-stationary stochastic bandit with a drift condition:
$\mathbb{P}(\overline{X}_{is}-\mu_{i}\geq\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})\>\text{and}\>\mathbb{P}(\overline{X}_{is}-\mu_{i}\leq-\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})$ (2)
Where $\overline{X}_{is}$ is the average reward of the first $s$ samples of arm $i$. $\mu_{i}$ is the limit of $\mathbb{E}[\overline{X}_{is}]$ as $s$ approaches infinity, which indicates that the expectation of the node value converges. $C$ is an appropriate constant characterizes the rate of concentration.
Based on this modeling, UCT uses the bandit algorithm UCB1 [8] to select child nodes. AlphaZero [9] inherits this concept and employs a variant formula:
@@ 4.1 Backward-view Reanalyze
Our algorithm design stems from a simple inspiration: if we could know the true state-value of a child node in advance, we could save the search for it, thus conserving search time. As shown on the right side of Figure 1, we directly use the true expected value $\mu_{A}$ to evaluate the quality of Arm A, thereby eliminating the need for the back-propagated value to calculate the score in 3. This results in the process occurring within the gray box being omitted. Indeed, this situation can be well modeled as an one-armed bandit[18], which also facilitates our subsequent theoretical analysis.
Figure 1: The connection between MCTS and bandits. Left shows the tree search in a bandit-view. When action $A$ is selected, a reward $r_{A}$ will be returned, where $r_{A}=\sum_{t=1}^{3}\gamma^{t-1}r_{t}^{a}+\gamma^{3}v_{a}$. For the root node, the traversal, evaluation and back-propagation occurring in the sub-tree can be approximated as sampling from an non-stationary distribution. Thus the root node can be seen as a non-stationary bandit. Right shows the one-armed bandit case. Once the reward $\mu_{A}$ is known, we can evaluate arm A using $\mu_{A}$, thereby eliminating the need to rely on subsequent tree search processes.
Driven by the aforementioned motivation, we aspire to obtain the reward expectation of a child node in advance. However, the true expectation is always unknown. Consequently, we resort to using the root value obtained from MCTS as an approximate substitute. Specifically, for two adjacent time steps $S_{0}^{1}$ and $S_{0}^{0}$  [^1] The subscript denotes the trajectory, the superscript denotes the time step in the trajectory. in Figure 2, when searching for state $S_{0}^{1}$, the root node corresponds to a child node of $S_{0}^{0}$. Therefore, the root value obtained can be utilized to assist the search for $S_{0}^{0}$. However, there is a temporal contradiction. $S_{0}^{1}$ is the successor state of $S_{0}^{0}$, yet we need to complete the search for $S_{0}^{1}$ first. Fortunately, this is possible during the reanalyze process.
Algorithm 1 shows the specific design with Python-like code. During reanalyze, since the trajectories was already collected in the replay buffer, we can perform tree search in a backward-view, which searches the trajectory in a reverse order. Figure 2 illustrates a batch containing $n+1$ trajectories, each of length $k+1$, and $S_{l}^{t}$ is the $t$-th state in the $l$-th trajectory. We first conduct a search for all $S^{k}$s, followed by a search for all $S^{k-1}$s, and so on. After we search on state $S_{l}^{t+1}$  [^2] To facilitate the explanation, we have omitted details such as the latent space and do not make a deliberate distinction between states and nodes., the root value $m_{l}^{t+1}$ is obtained. When engaging in search on $S_{l}^{t}$, we assign the value of $S_{l}^{t+1}$ to the fixed value $m_{l}^{t+1}$. During traverse in the tree, we select the action $a_{root}$ for root node $S_{l}^{t}$ with the following equation:
@@
where $a$ refers to the action associated with a child node, $a_{l}^{t}$ is the action corresponding to $S_{l}^{t+1}$, and $r_{l}^{t}$ signifies the reward predicted by the dynamic model. If an action distinct from $a_{l}^{t}$ is selected, the simulation continues its traversal in accordance with the original setting [10]. If action $a_{l}^{t}$ is selected, this simulation is terminated immediately. Since the time used to search for node $S^{t+1}_{l}$ is saved, this enhanced search process is faster than the original version.
@@
4.2 Theoretical Analysis
AlphaZero selects child nodes using Equation 3 and takes the final action based on the visit counts. In the previous section, we replace Equation 3 with Equation 5, which undoubtedly impacts the visit distribution of child nodes. Additionally, since we use the root value as an approximation to true expectation, the error between the two may also affect the search results. To demonstrate the reliability of our algorithm, in this section, we model the root node as an non-stationary bandit and prove that, as the number of total visit increases, the visit distribution gradually concentrates on the optimal arm. Specifically, we have the following theorem:
Theorem 1
For a non-stationary bandit that conforms to the assumptions of Equation 2, denote the total number of rounds as $n$, and the times a sub-optimal arm $i$ is selected in $n$ rounds as $T_{i}(n)$, then selecting actions according to Equation 5 can ensure that $\frac{{\mathbb{E}[T_{i}(n)]}}{n}\to 0$ as $n\to\infty$. Specifically, if we know the $n$ times sample mean $\hat{\mu}^{*}_{n}$ of the optimal arm in advance, then $\mathbb{E}[T_{i}(n)]$ for all sub-optimal arm $i$ satisfies
$\mathbb{E}[T_{i}(n)]\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}$ (6)
Otherwise, if we have the $n$ times sample mean $\hat{\mu}_{ln}$ of a sub-optimal arm $l$, then for arm $l$,
$\mathbb{E}[T_{l}(n)]\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (7)
and for other sub-optimal arms,
$\mathbb{E}[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (8)
where $\Delta_{i}$ is the optimal gap for arm $i$, $\varepsilon$ is a constant in $(0,\Delta_{i})$, and $C$ is the constant in Equation 2.
We provide the complete proof and draw similar conclusions for AlphaZero in Appendix A. Additionally, our method has a lower upper bound for $\mathbb{E}[T_{i}(n)]$. This implies that our algorithm may yield visit distribution more concentrated on the optimal arm. This is a potential worth exploring in future work, especially in offline scenarios [13] where reanalyze becomes the sole method for policy improvement.
@@ Appendix A Proof Materials
[…]
@@
, then it holds that $E[\kappa]\leq 1+\frac{2\sqrt{a}}{\varepsilon}+\frac{C^{2}}{\varepsilon^{2}}$.
@@
$E[\kappa]$ $\leq u+\sum_{t=\lceil u\rceil}^{n}\mathbb{P}(\hat{\mu}_{t}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon)$ (10)
$\leq u+\sum_{t=\lceil u\rceil}^{n}\exp(-\frac{t(\varepsilon-\sqrt{\frac{a}{t^{2}}})^{2}}{C^{2}})$ (11)
@@
$=1+u+\frac{C^{2}}{\varepsilon^{2}}=1+\frac{2\sqrt{a}}{\varepsilon}+\frac{C^{2}}{\varepsilon^{2}}$ (14)
Proof for Theorem 1:
@@
We first present the upper bound of $\mathbb{E}[T_{i}(n)]$ for using the Equation 3 in AlphaZero. A slight adjustment to this proof yields the conclusion of Theorem 1. Denote $T_{i}(k)$ as the times that arm $i$ has been chosen until time $k$, $\hat{\mu}_{is}$ as the average of the first $s$ samples of arm $i$, which satisfies the concentration assumption in Equation 2, and $\hat{\mu}_{i}(k)=\hat{\mu}_{iT_{i}(k)}$. Then we have:
@@
$\leq 1+\sum_{t=2}^{n}\sum_{s=1}^{t-1}\exp(-\frac{s(\varepsilon+P_{1}\frac{\sqrt{t-1}}{1+s})^{2}}{C^{2}})$ (21)
$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{C^{2}}\varepsilon P_{1}\sqrt{t-1})\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{C^{2}})$ (22)
$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{C^{2}}\varepsilon P_{1}\sqrt{t-1})\frac{C^{2}}{\varepsilon^{2}}$ (23)
$\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (24)
Notes: In 19, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00. Thus, we can discard it. If this assumption doesn’t hold, we can choose to accumulate $t$ starting from a larger $t_{0}$(which satisfies $P_{1}\sqrt{t_{0}-1}>\mu_{1}$ as mentioned in the article). Starting the summation from such a $t_{0}$ ensures all terms of $s=0$ can still be discarded, and all add terms that $t\leq t_{0}$ can be bounded to $1$. This only changes the constant term and won’t affect the growth rate of regret. From 21 to 22, we just need to expand the quadratic term and do some simple inequality scaling. From 22 to 23, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{C^{2}})$ is a geometric sequence and scale it to $\frac{C^{2}}{\varepsilon^{2}}$. From 23 to 24, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
for 17, we have
@@
282+$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}$ (29)
@@
$E[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (30)
The proof of theorem 1 can be obtained by making slight modifications. In case we have $\hat{\mu}^{*}_{n}=\frac{1}{n}\sum_{t=1}^{n}\hat{\mu}_{t}$ where $\hat{\mu}_{t}$ is sampled from the optimal arm in advance,
@@
$\leq n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (34)
$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}$ (35)
@@
$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (39)
$\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (40)
and the bound of $E[T_{i}(n)]$ for $i\neq l$ keeps unchanged. ∎